# Target Specific De Novo Design of Drug Candidate Molecules with Graph Transformer-based Generative Adversarial Networks

### v6

@@ 4.7 Performance Metrics
The second set of metrics concerns the physicochemical properties of de novo molecules: quantitative estimate of drug-likeness (QED) and synthetic accessibility (SA). QED computes the drug-like quality of a molecule using molecular descriptors such as molecular weight, lipophilicity, hydrogen bond donors and acceptors, polar surface area (PSA), and others. The SA score assesses the difficulty of synthesising de novo generated molecules by comparing them with labelled molecular building blocks. Higher values of QED and lower values of SA indicate better performance. Detailed descriptions of the metrics can be found in Polykovskiy et al. [43] and Landrum [68].
@@ 4.8 Molecule Docking
For the docking study, the crystal structure of the selected target protein, AKT1 (PDB code “4GV1”) [42], was prepared by Protein Preparation Wizard Madhavi [75] in the Schrödinger Suite 2022-4 [76] with the OPLS4 force field. Missing hydrogen atoms were added, and water molecules were removed. The physical condition of pH was set as 7.4 ± 1.0 for atom typing. The binding site of AKT1 was defined as Ala177, Lys179, Lys182, Ala212, Glu228, Ala230, Glu234, Glu278, Thr291, Asp292, by cross checking with the binding data published in the literature [42]. These findings were utilized for grid generation. Glide software [77] was used to find the best binding poses for each ligand. Van der Waals radius scaling factor was set to 1.0 and partial charge cut-off value was set to 0.25. The docking calculations were made in the Standard Precision Mode (GlideScore SP). Results were visualized via PyMOL [78].
4.9 Bioactivity Prediction Procedure
@@
Both real and de novo molecules were subjected to deep learning-based drug-target interaction (DTI) prediction against the selected protein (AKT1). We employed our previously developed DTI prediction system entitled DEEPScreen for this purpose. Briefly, DEEPScreen employs 2-D image-based structural Kakule representations (300-by-300 pixels) of compounds (created by RDKit, Draw.MolToFile function) as input and processes them via deep convolutional neural networks to classify them as active or inactive against the target of interest [56]. For this, we first trained a DEEPScreen model for the AKT1 protein using experimental bioactivity data available in ChEMBL (v30). The training data comprised 1338 active and 1666 inactive small molecules (activity threshold was selected as pChEMBL value: 7, equivalent to 100 nM concentration in terms of IC50). The high pChEMBL threshold was chosen to maximise the activity prediction confidence. We randomly split the compound dataset into train, validation and test folds (80%, 10% and 10% of the data, respectively). We augmented the dataset by adding transformed versions of each active and inactive molecule created by 10-degree rotations (the images were rotated around the centre of the image), making a total of 36 images for each input sample. This was done to render the model rotation invariant. Other types of image transformations, such as translation, were irrelevant to our case since the 2-D image generation function of RDKit standardises the molecular drawings. We optimised hyper-parameters of the model with respect to the classification metrics on the validation fold and measured the finalised performance of the model on the independent hold-out test fold. The test performance of the model was found to be precision: 0.91, recall: 0.92, F1-score: 0.92, and MCC: 0.85, which was considered highly satisfactory. Afterwards, the 2-D structural images of the real and de novo molecules were generated using the same parameters of RDKit Draw.MolToFile. The images were run on the trained AKT1 model in the prediction mode. Details regarding the DEEPScreen system and its training can be obtained from Rifaioglu et al. [56].
@@ 6 Code availability
The source code and ready-to-use trained models are available at the DrugGEN GitHub repo: https://github.com/HUBioDataLab/DrugGEN. DrugGEN is also available as an online tool with a graphical interface at https://huggingface.co/spaces/HUBioDataLab/DrugGEN, where users can generate de novo molecules by employing the desired DrugGEN model.
@@ 7 Acknowledgments
This project was supported by TUBITAK-BIDEB 2247-A National Leader Researchers Program under project number 120C123.
@@ 8 Author Information & Contributions
HAG: Heval Ataş Güvenilir (hevalatas@gmail.com),
AK: Altay Koyaş (altay.koyas@metu.edu.tr),
@@
TD conceptualised the study and designed the general methodology. EC and HAG prepared the datasets and handled the protein featurisation process. AS, AU, ASR and TD determined the technical details of the fundamental model architecture. AU and AS prepared the original codebase and designed and implemented initial models. AU and MGY designed, implemented, trained, tuned and evaluated numerous model variants and constructed the finalised DrugGEN models. DCK and AK selected the protein target by reviewing the literature. AO and EC conducted the molecular filtering operations and physics-based (docking and molecular dynamics) experiments. HC and AU analysed the de novo generated molecules in the context of deep learning-based drug-target interaction prediction. AU, EC, AO, EB and TD evaluated and discussed the findings. EC, AU, AS and TD visualised the results and prepared the figures in the manuscript. AU, EC, AS, HC, HAG, AO and TD wrote the manuscript. AU, EC, AS, MGY and TD prepared the repository. OB and MGY constructed the online tool. TD, AO, EB and ASR supervised the overall study. All authors approved the manuscript.

### v7

@@ 4.4 MLP Baseline Models
The second baseline replaces the transformer encoder blocks in DrugGEN with graph convolutional network (GCN) layers in both the generator and discriminator. For the generator, three consecutive GCN layers are applied to the annotation and adjacency matrices to update the node embeddings. These are then reshaped into the generated molecule representation via a final readout layer, ensuring the correct output dimensions. In the discriminator, three GCN layers similarly encode the input molecule (real or generated) into node embeddings, which are then aggregated into a single feature vector. This vector is passed to a four-layer MLP (with an input size of 64, followed by hidden layers of 32 and 16 neurons, and a single output neuron), same with the DrugGEN discriminator, to produce the final scalar score.
We developed both MLP and GCN baselines (targeted and non-targeted versions) under the same training, hyperparameter, and evaluation regimes as DrugGEN to ensure a fair comparison. The results obtained from these baseline models are discussed in the ablation study.
@@ 4.7 Performance Metrics
The second set of metrics concerns the physicochemical properties of de novo molecules: quantitative estimate of drug-likeness (QED) and synthetic accessibility (SA). QED computes the drug-like quality of a molecule using molecular descriptors such as molecular weight, lipophilicity, hydrogen bond donors and acceptors, polar surface area (PSA), and others. The SA score assesses the difficulty of synthesising de novo generated molecules by comparing them with labelled molecular building blocks. Higher values of QED and lower values of SA indicate better performance. In addition to these metrics, the drug-likeness of molecules is further evaluated using Lipinski’s Rule of 5 [82], Veber’s rules [83], and the Pan-Assay Interference Compounds (PAINS) filter [84]. Lipinski’s Rule and Veber’s rules focus on optimising pharmacokinetics, bioavailability, and membrane permeability, while the PAINS filter eliminates compounds prone to non-specific binding or false positives, ensuring the quality of biological screening. Together, these rules and filters enhance the assessment of a molecule’s drug-like potential. We also included the Frechet ChemNet Distance (FCD) metric to assess the physicochemical similarity between molecular datasets. This metric uses the Frechet distance applied to molecular structure embeddings generated by the ChemNet model. Finally, to evaluate structural similarity, we employed scaffold and fragment similarity metrics from MOSES, which utilise Murcko scaffolds and BRICS decomposition for comparative analysis. Detailed descriptions of the metrics can be found in Polykovskiy et al. [49] and Landrum [75].
@@ 4.8 Molecule Docking
For the docking study, the crystal structure of the selected target protein, AKT1 (PDB ID: 4GV1) [48], was prepared by Protein Preparation Wizard Madhavi [85] in the Schrödinger Suite 2022-4 [86] with the OPLS4 force field. Missing hydrogen atoms were added, and water molecules were removed. The pH value was set to 7.4 ± 1.0 for atom typing. The binding site of AKT1 was defined based on the literature data and included Ala177, Lys179, Lys182, Ala212, Glu228, Ala230, Glu234, Glu278, Thr291, and Asp292 [48]. Docking studies were performed using Glide [87]to determine the top-scoring binding poses for each ligand. The Van der Waals radius scaling factor was set to 1.0, and the partial charge cut-off value was set to 0.25. Docking calculations were conducted in the Standard Precision (SP) mode using the GlideScore scoring function. The same docking procedure was applied for CDK2 (PDB ID: 4KD1) [88]. The crystal structure of CDK2 was prepared similarly, and its binding site was confirmed based on literature data [88]. All results were visualized via PyMOL [89].
4.9 Dimensionality reduction with UMAP and t-SNE
Figure 2B present the UMAP (parameters: $n_{n}eighbors$=50, $min_{d}ist$=0.8 and metric=”dice”) and t-SNE (parameters: perplexity: 500, iteration: 2000) projections, respectively, of a randomly selected subset of de novo molecules from DrugGEN and DrugGEN-NoTarget models (1,000 from each), real AKT1 inhibitors (all molecules used in model training); and 50,000 ChEMBL molecules from the training dataset. In these plots, each dot represents a molecule, colours indicate their source, and the Euclidean distances roughly indicate structural similarities based on Tanimoto applied on molecular fingerprints using the MACCS (Molecular ACCess System) descriptors [90].
4.10 DEEPScreen Model Training and Evaluation Procedure
Both real and de novo molecules were subjected to deep learning-based drug-target interaction (DTI) prediction against the selected protein (AKT1). We employed our previously developed DTI prediction system, entitled DEEPScreen, for this purpose. Briefly, DEEPScreen employs 2-D image-based structural Kakule representations (300-by-300 pixels) of compounds (created by RDKit, Draw.MolToFile function) as input and processes them via deep convolutional neural networks to classify them as active or inactive against the target of interest [57]. We preferred DEEPScreen for this analysis since (i) it is straightforward to prepare input data and train/test a target-specific DTI prediction model, (ii) it only requires bioactivity data, (iii) it has high predictive performance [57]. DEEPScreen is completely separate/independent from DrugGEN in terms of the modelling approach, datasets, and output. Therefore, this DTI prediction-based validation is objective.
@@ 4.10 DEEPScreen Model Training and Evaluation Procedure
For this, we first trained a DEEPScreen model for the AKT1 protein using experimental bioactivity data available in ChEMBL (v35). The training data comprised 1932 active and 2414 inactive small molecules (activity threshold was selected as pChEMBL value: 6, equivalent to 1 µM concentration in terms of IC50). We split the compound dataset into train, validation and test folds using the balanced scaffold split function from the chemprop library [91]. By ensuring a balanced label ratio across the folds the dataset is partitioned into training, validation, and test sets containing 3.090, 774, and 968 molecules, respectively. We augmented the dataset by adding transformed versions of each active and inactive molecule created by 10-degree rotations (the images were rotated around the centre of the image), making a total of 36 images for each input sample. This was done to render the model rotation invariant. Other types of image transformations, such as translation, were irrelevant to our case since the 2-D image generation function of RDKit standardises the molecular drawings. We evaluated the model with a confidence score threshold of 0.5 for the binary classification task (active/inactive), classifying molecules with scores $\geq$ 0.5 as active which corresponds to $\geq$ 18 augmented images. We optimised the hyper-parameters of the model with respect to the classification metrics on the validation fold and measured the finalised performance of the model on the independent hold-out test fold. Afterwards, the 2-D structural images of the real and de novo molecules were generated using the same parameters of RDKit Draw.MolToFile. The images were run on the trained AKT1 model in the prediction/inference mode. Details regarding the DEEPScreen system and its training can be obtained from https://github.com/HUBioDataLab/DEEPScreen2.
@@ 4.11 Molecular Dynamics Simulations
4.12 Experimental procedures for enzymatic assays
Our experimental validation analyses were designed to assess the direct binding of de novo generated molecules to AKT1, leading to the specific inhibition of the enzyme. To accomplish this, we outsourced the enzymatic assays to Reaction Biology Europe GmbH in Germany. Using Reaction Biology’s radiometric 33PanQinaseTM activity assay, we obtained the IC50 profiles of the synthesised compounds against AKT1. The detailed assay methodology can be found in the Supplementary Material (S12).
@@ 6 Code availability
The source code and ready-to-use trained models are available in the archived DrugGEN repository (DOI:10.5281/zenodo.15014579) and can be accessed at https://github.com/HUBioDataLab/DrugGEN. DrugGEN is also available as an online tool with a graphical interface at https://huggingface.co/spaces/HUBioDataLab/DrugGEN, where users can generate de novo molecules by employing the desired model.
@@ 7 Acknowledgments
This project was supported by TUBITAK-BIDEB 2247-A National Outstanding Researchers Program under project number 120C123. The authors thank Heval Ataş Güvenilir for guidance during the preparation of datasets and Altay Koyaş for aiding the target protein selection process.
@@ 8 Author Information & Contributions
TD conceptualised the study and designed the general methodology. EC prepared the datasets. AS, AU, ASR and TD determined the technical details of the fundamental model architecture. AU and AS prepared the original codebase and designed and implemented initial models. AU and MGY designed, implemented, trained, tuned and evaluated numerous model variants and constructed the finalised DrugGEN models. AO and EC conducted the molecular filtering operations and physics-based (docking and molecular dynamics) experiments. AU and HC analysed the de novo generated molecules in the context of deep learning-based drug-target interaction prediction. MGY, EC and TD conducted the attention map analysis. AU, EC, AO, EB and TD evaluated and discussed the findings. EC, AU, AS, AO and TD visualised the results and prepared the figures in the manuscript. AU, EC, MGY, AS, DCK, AO and TD wrote the manuscript. AU, EC, AS, MGY and TD prepared the repository. OB and MGY constructed the online tool. TD supervised the overall study. All authors approved the manuscript.